For integers 𝑥 and 𝑦, consider the proposition ∃𝑥∀𝑦 (𝑥 + 𝑦 = 10). Is the proposition true, false, or does it depend on the values of 𝑥 and 𝑦?


1
Expert's answer
2021-10-14T05:25:55-0400

The proposition "\u2203\ud835\udc65\u2200\ud835\udc66 (\ud835\udc65 + \ud835\udc66 = 10)" is false because of for any integer number "x" there exists the integer number "y=-x" such that "x+y=x-x=0\ne 10."